How do I solve an unlike rational expression?
I need to solve:
9/x-5 minus 1=8/x+5
here is what I have so far
(x+5)(x-5)9/x-5 minus 1(x+5)(x-5)=8/x+5 (x+5)(x-5)
this gives me 7x+45=8x-40 so x=85
If I try and plug this back into the equation, it doesn't seem to equate properly. what did I do wrong? Thanks!

Answer by stanbon(48569)

9/x-5 minus 1=8/x+5
here is what I have so far
(x+5)(x-5)9/x-5 minus 1(x+5)(x-5)=8/x+5 (x+5)(x-5)
Cancel where you can to get:
9(x+5) - (x^2-25) = 8(x-5)
9x+45 - x^2+25 = 8x-40
x^2-x-110=0
(x-11)(x+10)=30
x=11 or x=-10
Cheers,
Stan H.